# Limit-case admissibility for positive infinite-dimensional systems

### v1

@@ 3.2 Zero-class $L^{1}$-admissibility
The assertion is now immediate.
A subset $S$ of an ordered Banach space $X$ is called almost order-bounded if for each $\epsilon>0$ there exists $x_{\epsilon}\in X_{+}$ such that $\left\lVert(\left\lvert x\right\rvert-x_{\epsilon})_{+}\right\rVert<\epsilon$ for all $x\in S$. Equivalently, if for each $\epsilon>0$ there exists $x_{\epsilon}\in X_{+}$ such that
$S\subseteq[-x_{\epsilon},x_{\epsilon}]+\epsilon B_{X};$
where $B_{X}$ denotes the unit ball of $X$. Moreover, an AL-space is a Banach lattice whose norm is additive on the positive cone (cf. Table 1).
@@
Proof 3.14.
Let $(R_{t})_{t\geq 0}$ denote the nilpotent right-translation semigroup on the space $L^{1}([0,1],Y)$. Reflexivity of $X$ implies weak compactness of the output operator $\Psi_{1}$. In turn, $\Psi_{1}(B_{X})$ is relatively weakly compact. Next, as $Y$ is an AL-space, so is $L^{1}([0,1],Y)$. Consequently, $\Psi_{1}(B_{X})$ is almost order-bounded by Dunford-Pettis theorem [38, Theorem 2.5.4]. Strong continuity and nilpotency of $(R_{t})_{t\geq 0}$ now implies that it converges to 00 uniformly on $\Psi_{1}(B_{X})$ as $t\uparrow 1$. Consequently, $C$ is zero-class $L^{1}$-admissible by Lemma 3.11.
Note that all of the order assumptions in Theorem 3.13 are posed on $Y$ whereas $X$ is a general Banach space. We don’t know whether the Dunford-Pettis theorem [38, Theorem 2.5.4] can be generalised to ordered Banach spaces whose norm is additive on the positive cone. For this reason, the proof above necessitated that $Y$ be a Banach lattice.
@@
In the proof of the above theorem, reflexivity is only used to obtain weak compactness of $\Psi_{1}$. In other words, for $L^{1}$-admissible observation operators mapping into an AL-space, weak compactness of the output operator implies zero-class $L^{1}$-admissibility.
The following example shows that the reflexivity assumption in Theorem 3.13 cannot be dropped, in general; see also [33, Example 26].
@@
Consider the left translation semigroup on $X:=L^{1}[0,1]$ with observation $C:X_{1}\to\mathbb{C}$ given by $f\mapsto f(0)$. The corresponding input operator $\Psi_{\tau}:X_{1}\to L^{1}([0,\tau],Y)$ satisfies $\left\lVert\Psi_{\tau}\right\rVert=1$ for each $\tau>0$. Therefore, $C$ is $L^{1}$-admissible but the admissibility is not zero-class.
Combining Proposition 3.1 with Theorem 3.13 yields the following sufficient condition for zero-class $L^{1}$-admissibility of positive observation operators:
@@
Suppose that Assumption 3 holds and assume that $X$ is reflexive and $Y$ is an AL-space. If $C\in\mathcal{L}(X_{1},Y)$ is positive, then it is zero-class $L^{1}$-admissible.
Yet another condition for zero-class $L^{1}$-admissibility of observation operators is given in Lemma 5.5.
4 Admissibility of control operators
In this section, we shift our attention to the admissibility considerations of the system
@@ 4 Admissibility of control operators
We make the following assumption on the $C_{0}$-semigroup $(T(t))_{t\geq 0}$ generated by $A$ and the control operator $B\in\mathcal{L}(U,X_{-1})$:
{assumption}
Suppose that $X$ is an ordered Banach space with a generating and normal cone and $(T(t))_{t\geq 0}$ is a positive $C_{0}$-semigroup on $X$. Moreover, $U$ is a Banach space and $B\in\mathcal{L}(U,X_{-1})$.
In what follows, we let $\mathrm{T}([0,\tau],U)$ denote the space of all $U$-valued step functions on $[0,\tau]$, i.e., the functions that can be written as a finite linear combination of characteristic functions of intervals. This is a normed space when equipped with the supremum norm whose closure is called the space of regulated functions – denoted by $\operatorname{Reg}([0,\tau],U)$.
Now, let $\mathrm{Z}$ be a placeholder for $L^{\infty}$ or $\mathrm{C}$ or $\operatorname{Reg}$. Corresponding to the system in (4.1), we define the input map as
(4.2) $\Phi_{\tau}:\mathrm{Z}([0,\tau],U)\to X_{-1},\qquad u\mapsto\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s$
for fixed $\tau>0$. The operator $B$ is called $\mathrm{Z}$-admissible control operator if $\operatorname{Rg}\Phi_{\tau}\subseteq X$ for some (equivalently, all) $\tau>0$. Additionally, if $\limsup_{\tau\downarrow 0}\left\lVert\Phi_{\tau}\right\rVert_{\mathcal{L}(\mathrm{Z}([0,\tau],U),X)}=0$, then we say that $B$ is a zero-class $\mathrm{Z}$-admissible control operator.
Since the space of regulated functions contains the space of continuous functions, so every $\operatorname{Reg}$-admissible control operator is $\mathrm{C}$-admissible. In fact, the converse is also true and seems to be folklore. The proof for the particular case $U=X$ and $B=A_{-1}$ is given in [32, Proposition 2.2] and the proof for the general case follows mutatis mutandis. As we use this observation multiple times in the sequel, we state it explicitly:
@@
4.1 Positivity of $B$ in terms of the boundary control operator
As explained in the introduction, the additive input term $Bu$ in (4.1) with an operator $B:U\to X_{-1}$ is often used to encode the boundary control. Since several results in this section use the assumption that $B$ is positive, we will first describe how the positivity of $B$ can be described in terms of the boundary operator. The precise setting is as follows.
Let $X$ and $U$ be ordered Banach spaces and let $\mathfrak{A}:\operatorname{dom}\mathfrak{A}\subset X\to X$ and $\mathfrak{B}:\operatorname{dom}\mathfrak{A}\to U$ be linear operators. Assume that $\mathfrak{B}$ has a bounded right-inverse and that the restriction $A:=\mathfrak{A}|_{\ker\mathfrak{B}}$ generates a positive $C_{0}$-semigroup on $X$. For every $\lambda\in\rho(A)$, one has the following properties (see for instance [24, Lemma 1.2]): the domain $\operatorname{dom}\mathfrak{A}$ is the direct sum of its vector subspaces $\operatorname{dom}(A)=\ker\mathfrak{A}$ and $\ker(\lambda-\mathfrak{A})$ and the restriction $\mathfrak{B}|_{\ker(\lambda-\mathfrak{A})}$ is a bijection from $\ker(\lambda-\mathfrak{A})$ to $U$. We denote its inverse by $B_{\lambda}:U\to\ker(\lambda-\mathfrak{A})\subseteq\operatorname{dom}\mathfrak{A}\subseteq X$. The boundary control problem
@@ 4.1 Positivity of $B$ in terms of the boundary control operator
with control $u(t)\in U$ can, under appropriate assumptions, be reformulated as the system (4.1) if one chooses $B:=(\lambda-A_{-1})B_{\lambda}:U\to X_{-1}$, where $A_{-1}$ is the generator of the extrapolated $C_{0}$-semigroup on $X_{-1}$; see [48, Proposition 2.8]. One can check that the operator $B$ does not depend on the choice of $\lambda$ (see for instance [24, Formula (1.16) in Lemma 1.3]). We now show how positivity of $B$ can be characterised in terms of the operators $B_{\lambda}$.
@@
In the setting described before the proposition, the following assertions are equivalent:
@@
Proof 4.3.
“(i) $\Rightarrow$ (iii)”: Assume that (i) holds and let $\lambda>s(A)$. Then $R(\lambda,A_{-1})$ is a positive operator from $X_{-1}$ to $X$ according to Proposition 2.12(c). Hence, $B_{\lambda}=R(\lambda,A_{-1})B$ is positive.
@@
where both limits are norm limits in $X_{-1}$.
We point out that the equivalence of (i) and (iii) in Proposition 4.2 above was also noted in [19, Lemma 2.1] and on [20, Page 16].
In a similar vein, if one already knows that a boundary control system is admissible, then the positivity of the operators $B_{\lambda}$ characterises whether positive inputs lead to positive trajectories [16, Proposition 4.3]. The condition that all $B_{\lambda}$ be positive also occurs in assumption (ii) of [19, Theorem 4.3]. It is instructive to discuss what Proposition 4.2 says for the Laplace operator. The following example demonstrates that the positivity condition on $B_{\lambda}$ can be interpreted as the maximum principle for the Laplace operator in this case.
@@
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded smooth domain (say, for the sake of simplicity, with $C^{\infty}$-boundary), set $X:=L^{p}(\Omega)$ for some $p\in(1,\infty)$ and let $\mathfrak{A}$ denote the Laplace operator with domain $W^{2,p}(\Omega)$. Let $U\subseteq L^{p}(\partial\Omega)$ denote the image of the trace operator on $W^{2,p}(\Omega)$, which we endow with the order inherited from $L^{p}(\partial\Omega)$, and let $\mathfrak{B}:W^{2,p}(\Omega)\to U$ be the trace operator.
Then $A$ is the Dirichlet Laplace operator on $L^{p}(\Omega)$. For $\lambda>s(A)$ and $u\in U$, the function $v:=B_{\lambda}u\in W^{2,p}(\Omega)$ solves the boundary value problem
@@
So by the maximum principle, the operator $B_{\lambda}$ is positive. Hence, Proposition 4.2 yields that $B$ is positive.
4.2 Input admissibility for positive semigroups
If $U$ and $X$ in Assumption 4 are reflexive and if $B\in\mathcal{L}(U,X_{-1})$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$, then the same assumption holds for the double dual $B^{\prime\prime}\in\mathcal{L}(U^{\prime\prime},(X^{\prime\prime})_{-1})$. Due to the reflexivity of $U$, we obtain that $B^{\prime}$ is an $L^{1}$-admissible observation operator by Corollary 3.9. Therefore, reflexivity of $X$ allows us to appeal to the Weiss duality theorem [55, Theorem 6.9(ii)] to obtain that $B$ is $L^{\infty}$-admissible. Without the reflexivity assumptions on $U$ and $X$, we are able to show that $L^{\infty}$-admissibility is always zero-class.
@@ 4.2 Input admissibility for positive semigroups
Suppose that Assumption 4 holds and that $B\in\mathcal{L}(U,X_{-1})$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$. If $B$ is $L^{\infty}$-admissible, then it is zero-class $L^{\infty}$-admissible.
In particular, $B$ is always zero-class $L^{\infty}$-admissible if both $X$ and $U$ are reflexive.
Proof 4.6.
By assumption, there exist $a,b\in X_{-1}$ such that $B(B_{U})\subseteq[a,b]$. Without loss of generality, assume that $\omega_{0}(A)<0$ and that the norm on $X_{-1}$ is given by $\left\lVert x\right\rVert_{-1}=\left\lVert(A_{-1})^{-1}x\right\rVert_{X}$ for all $x\in X_{-1}$. Suppose that $B$ is $L^{\infty}$-admissible. Then the corresponding input operator $\Phi_{\tau}:L^{\infty}([0,\tau],U)\to X_{-1}$ maps into $X$ and for each $u\in L^{\infty}([0,\tau],U)$ taking values in $B_{U}$, we get
$\left\lVert\Phi_{\tau}u\right\rVert_{X}\leq\left\lVert\int_{0}^{\tau}T_{-1}(s)b\;\mathrm{d}s\right\rVert_{X}=\left\lVert A_{-1}\int_{0}^{\tau}T_{-1}(s)b\;\mathrm{d}s\right\rVert_{-1}=\left\lVert T_{-1}(\tau)b-b\right\rVert_{-1}\to 0$
as $\tau\to 0$. In other words, $B$ is zero-class $L^{\infty}$-admissible. The second assertion can now be inferred from the discussion preceding the theorem.
Remark 4.7.
With the same arguments as above, it can be shown that the first assertion of Theorem 4.5 remains true if $L^{\infty}$ is replaced with $\mathrm{C}$.
@@
Suppose that Assumption 4 holds and that $B\in\mathcal{L}(U,X_{-1})$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$. Then $B$ is zero-class $C$-admissible.
Proof 4.9.
For each $u\in\mathrm{T}([0,\tau],U)$ – the space of all $U$-valued step functions on $[0,\tau]$ – we know that
@@
from [10, Lemma 4.3(ii)]; note that the proof of [10, Lemma 4.3(ii)] doesn’t use any positivity properties. The positivity of the semigroup and the assumption on $B$ guarantees that there exists $b\in X_{-1}$ such that if $u\in\mathrm{T}([0,\tau],U)$ with $\left\lVert u\right\rVert_{\infty}\leq 1$, then
$\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s\right\rVert\leq\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)b\;\mathrm{d}s\right\rVert=:K_{\tau}$
@@
for all step functions $u\in\mathrm{T}([0,\tau],U)$. Hence, $B$ is $\mathrm{C}$-admissible (see, [56, Remark 4.1.7] and Proposition 4.1). The zero-class $\mathrm{C}$-admissibility now follows from Remark 4.7.
In particular, the following generalisation of [56, Corollary 4.2.10] follows by combining Theorem 4.8 and Proposition A.3(a).
@@
Suppose that Assumption 4 holds and let $U$ be an ordered Banach space with a unit. Then every positive $B\in\mathcal{L}(U,X_{-1})$ is zero-class $\mathrm{C}$-admissible.
Another situation where $B$ satisfies the assumptions of Theorem 4.8 is given in Proposition A.3(b). However, such a condition even yields zero-class $L^{\infty}$-admissibility as we see in Corollary 4.16.
@@
The assumption that $U$ has a unit cannot be dropped in Corollary 4.10. Indeed, on the Banach lattice $c_{0}$, there is an example [32, Example 2.3] – due to Kato – of a positive semigroup whose generator is positive, $\mathrm{C}$-admissible, yet it is not zero-class $\mathrm{C}$-admissible.
Recall from Table 1 that an AM-space is a Banach lattice whose closed unit ball is upwards directed. In particular, $c_{0}$ is an AM-space. It was shown in [56, Theorem 4.1.19] that if $U$ is an AM-space, then every positive $B\in\mathcal{L}(U,X_{-1})$ is at least $\mathrm{C}$-admissible. We generalise this below:

### v4

@@ 3.2. Zero-class $L^{1}$-admissibility
The assertion is now immediate. ∎
A subset $S$ of a Banach lattice $X$ is called almost order-bounded if for each $\varepsilon>0$ there exists $x_{\varepsilon}\in X_{+}$ such that $\left\lVert(\left\lvert x\right\rvert-x_{\varepsilon})_{+}\right\rVert<\varepsilon$ for all $x\in S$. Equivalently, if for each $\varepsilon>0$ there exists $x_{\varepsilon}\in X_{+}$ such that
$S\subseteq[-x_{\varepsilon},x_{\varepsilon}]+\varepsilon B_{X};$
where $B_{X}$ denotes the closed unit ball of $X$. Recall that an AL-space is a Banach lattice whose norm is additive on the positive cone (cf. Table 1).
@@
Proof.
Reflexivity of $X$ implies weak compactness of the output operator $\Psi_{1}$, so $\Psi_{1}(B_{X})$ is relatively weakly compact. Next, as $Y$ is an AL-space, so is $L^{1}([0,1],Y)$. Consequently, $\Psi_{1}(B_{X})$ is almost order-bounded by the Dunford-Pettis theorem [44, Theorem 2.5.4]. Letting $(R_{t})_{t\geq 0}$ be as in Lemma 3.8, its strong continuity and nilpotency now implies that it converges to 00 uniformly on $\Psi_{1}(B_{X})$ as $t\uparrow 1$. Consequently, $C$ is zero-class $L^{1}$-admissible by Lemma 3.8. ∎
Note that all of the order assumptions in Theorem 3.9 are posed on $Y$ whereas $X$ is a general Banach space. We don’t know if the Dunford-Pettis theorem [44, Theorem 2.5.4] can be generalised to ordered Banach spaces whose norm are additive on the positive cone. Therefore, the proof above required that $Y$ is a Banach lattice.
@@
In the proof of Theorem 3.9, reflexivity was needed solely for $\Psi_{1}$ to be weakly compact. In other words, for AL-space valued $L^{1}$-admissible observation operators, weak compactness of the output map implies zero-class $L^{1}$-admissibility.
The following example shows that the reflexivity assumption in Theorem 3.9 cannot be dropped, in general; see also [38, Example 26].
@@
Consider the left translation semigroup on $X:=L^{1}[0,1]$ with observation $C:X_{1}\to\mathbb{C}$ given by $f\mapsto f(0)$. The corresponding output operator $\Psi_{\tau}:X_{1}\to L^{1}([0,\tau],Y)$ satisfies $\left\lVert\Psi_{\tau}\right\rVert=1$ for each $\tau>0$. Therefore, $C$ is $L^{1}$-admissible but the admissibility is not zero-class.
Combining Proposition 3.2 with Theorem 3.9 yields the following sufficient condition for zero-class $L^{1}$-admissibility of positive observation operators. Yet another condition for zero-class $L^{1}$-admissibility of observation operators is given in Lemma 5.4.
@@
Suppose that Assumption 3.1 holds. If $X$ is reflexive, $Y$ is an AL-space and $C\in\mathcal{L}(X_{1},Y)$ is positive, then $C$ is zero-class $L^{1}$-admissible.
4. Admissibility of control operators
In this section, we shift our attention to admissibility considerations regarding the system
@@ 4. Admissibility of control operators
where $A$ generates a $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on a Banach space $X$ and the control operator $B\in\mathcal{L}(U,X_{-1})$ is defined on a Banach space $U$. Most of our results in this section are proved in the following setting:
Assumption 4.1.
Suppose that $X$ is an ordered Banach space with a generating and normal cone and $(T(t))_{t\geq 0}$ is a positive $C_{0}$-semigroup on $X$. Moreover, let $U$ be a Banach space and $B\in\mathcal{L}(U,X_{-1})$.
In what follows, we let $\mathrm{T}([0,\tau],U)$ denote the space of all $U$-valued step functions on $[0,\tau]$, i.e., a piecewise constant function with finitely many jumps. This is a normed space when equipped with the supremum norm whose closure is called the space of regulated functions – denoted by $\operatorname{Reg}([0,\tau],U)$. Now, let $\mathrm{Z}$ be a placeholder for $L^{\infty}$ or $\mathrm{C}$ or $\operatorname{Reg}$. Corresponding to the system in (4.1), we define the input map
$\Phi_{\tau}:\mathrm{Z}([0,\tau],U)\to X_{-1},\qquad u\mapsto\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s$ (4.2)
for fixed $\tau>0$. The operator $B$ is called a $\mathrm{Z}$-admissible control operator if $\operatorname{Rg}\Phi_{\tau}\subseteq X$ for some (equivalently, all) $\tau>0$. Additionally, if $\limsup_{\tau\downarrow 0}\left\lVert\Phi_{\tau}\right\rVert_{\mathcal{L}(\mathrm{Z}([0,\tau],U),X)}=0$, then we say that $B$ is a zero-class $\mathrm{Z}$-admissible control operator. The admissibility of control operators was recently generalized to a notion closely related to maximal regularity in [11, Section 2]
For $\omega\in\mathbb{C}$, it is easy to see that $B$ is $\mathrm{Z}$-admissible for $(T(t))_{t\geq 0}$ if and only if $B$ is $\mathrm{Z}$-admissible for $(e^{\omega t}T(t))_{t\geq 0}$. For this reason, when proving admissibility of control operators, we often assume (without loss of generality) that $\omega_{0}(A)<0$ thereby having $0\in\rho(A)$.
Since $C([0,\tau],U)\subseteq\operatorname{Reg}([0,\tau],U)$, every $\operatorname{Reg}$-admissible control operator is $\mathrm{C}$-admissible. In fact, the converse is also true and seems to be folklore. The proof for the particular case $U=X$ and $B=A_{-1}$ is given in [37, Proposition 2.2] and the proof for the general case follows mutatis mutandis; see also [12, Proposition 2.1]. As we use this observation multiple times in the sequel, we state it explicitly:
@@
4.1. Positivity of $B$ in terms of the boundary control operator
As explained in the introduction, the additive input term $Bu$ in (4.1) with an operator $B:U\to X_{-1}$ is often used to encode boundary control. Since several results in this section use the assumption that $B$ is positive, we will first describe how the positivity of $B$ can be described in terms of the boundary operator. The precise setting is as follows.
Let $X$ and $U$ be ordered Banach spaces, let $\operatorname{dom}\mathfrak{A}$ be a vector subspace of $X$ that is a Banach space with respect to a stronger norm, and let $\mathfrak{A}:\operatorname{dom}\mathfrak{A}\to X$ and $\mathfrak{B}:\operatorname{dom}\mathfrak{A}\to U$ be bounded linear operators (but the norm on $\operatorname{dom}\mathfrak{A}$ does not need to be the graph norm of $\mathfrak{A}$ nor does $\mathfrak{A}$ need to be closed as an operator on $X$). Assume that $\mathfrak{B}$ is surjective and that the restriction $A:=\mathfrak{A}|_{\ker\mathfrak{B}}$ generates a positive $C_{0}$-semigroup on $X$. For every $\lambda\in\rho(A)$, one has the following properties (see e.g. [29, Lemma 1.2]): the domain $\operatorname{dom}\mathfrak{A}$ is the direct sum of its subspaces $\operatorname{dom}(A)=\ker\mathfrak{B}$ and $\ker(\lambda-\mathfrak{A})$ and the restriction $\mathfrak{B}|_{\ker(\lambda-\mathfrak{A})}$ is a bijection from $\ker(\lambda-\mathfrak{A})$ to $U$. We denote its inverse by $B_{\lambda}:U\to\ker(\lambda-\mathfrak{A})\subseteq\operatorname{dom}\mathfrak{A}\subseteq X$ and note that $B_{\lambda}$ is bounded from $U$ to $\operatorname{dom}\mathfrak{A}$ by the continuous inverse theorem. The boundary control problem
@@ 4.1. Positivity of $B$ in terms of the boundary control operator
with control $u(t)\in U$ can, under appropriate assumptions, be reformulated as the system (4.1) if one chooses $B:=(\lambda-A_{-1})B_{\lambda}:U\to X_{-1}$, where $A_{-1}$ is the generator of the extrapolated $C_{0}$-semigroup on $X_{-1}$; see [54, Proposition 2.8]. One can check that the operator $B$ does not depend on the choice of $\lambda$ (see for instance [29, Formula (1.16) in Lemma 1.3]). We now characterise positivity of $B$ in terms of $B_{\lambda}$.
@@
In the setting above, the following assertions are equivalent:
@@
Proof.
“(i) $\Rightarrow$ (iii)”: Fix $\lambda>s(A)$. Then $R(\lambda,A_{-1})$ is a positive operator from $X_{-1}$ to $X$ according to Proposition 2.8(c). Hence, $B_{\lambda}=R(\lambda,A_{-1})B$ is positive.
@@
where both limits are norm limits in $X_{-1}$. ∎
We point out that the equivalence of (i) and (iii) in Proposition 4.3 above was also noted in [22, Lemma 2.1] and on [23, Page 16]. In a similar vein, if one already knows that a boundary control system is admissible, then the positivity of the operators $B_{\lambda}$ characterises whether positive inputs lead to positive trajectories [19, Proposition 4.3]. The condition that all $B_{\lambda}$ be positive also occurs in assumption (ii) of [22, Theorem 4.3]. It is instructive to see what Proposition 4.3 says for the Laplace operator. The following example demonstrates that the positivity condition on $B_{\lambda}$ can be interpreted as the maximum principle for the Laplace operator in this case.
@@
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded smooth domain (say, for simplicity, with $C^{\infty}$-boundary), set $X:=L^{p}(\Omega)$ for some $p\in(1,\infty)$, and let $\mathfrak{A}$ denote the Laplace operator with domain $W^{2,p}(\Omega)$. Let $U\subseteq L^{p}(\partial\Omega)$ denote the image of the trace operator on $W^{2,p}(\Omega)$, endowed with the order inherited from $L^{p}(\partial\Omega)$ and let $\mathfrak{B}:W^{2,p}(\Omega)\to U$ be the trace operator.
Then, in the notation of Proposition 4.3, $A$ is the Dirichlet Laplace operator on $L^{p}(\Omega)$. For $\lambda>s(A)$ and $u\in U$, the function $v:=B_{\lambda}u\in W^{2,p}(\Omega)$ solves the boundary value problem
@@
So if $p\geq 2$, then it follows from the maximum principle for Sobolev functions [24, Theorem 8.1] that the operator $B_{\lambda}$ is positive. Hence, Proposition 4.3 yields that $B$ is positive if $p\geq 2$.
4.2. Input admissibility for positive semigroups
If $U$ and $X$ in Assumption 4.1 are reflexive and if $B\in\mathcal{L}(U,X_{-1})$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$, then the same assumption holds for the double dual $B^{\prime\prime}\in\mathcal{L}(U^{\prime\prime},(X^{\prime\prime})_{-1})$. Due to the reflexivity of $U$, we obtain that $B^{\prime}$ is an $L^{1}$-admissible observation operator by Corollary 3.7. Therefore, reflexivity of $X$ allows us to appeal to the Weiss duality theorem [61, Theorem 6.9(ii)] to obtain that $B$ is $L^{\infty}$-admissible. Without the reflexivity assumptions on $U$ and $X$, we are able to show that $L^{\infty}$-admissibility is always zero-class.
In what follows, we repeatedly use the following property of normal cones. If $X$ is an ordered Banach space with a normal cone, then by [4, Theorem 2.38] there exists $c>0$ such that
$x\in[a,b]\quad\Rightarrow\quad\left\lVert x\right\rVert\leq c\max\{\left\lVert a\right\rVert,\left\lVert b\right\rVert\}.$ (4.3)
@@ 4.2. Input admissibility for positive semigroups
Suppose that Assumption 4.1 holds and that $B\in\mathcal{L}(U,X_{-1})$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$.
[…]
    If $B$ is $L^{\infty}$-admissible, then it is zero-class $L^{\infty}$-admissible.
[…]
    If both $X$ and $U$ are reflexive, then $B$ is zero-class $L^{\infty}$-admissible.
[…]
    If $B$ is $\mathrm{C}$-admissible, then it is zero-class $\mathrm{C}$-admissible.
Proof.
(a) By assumption, there exist $b_{1},b_{2}\in X_{-1}$ such that $B(B_{U})\subseteq[b_{1},b_{2}]$. Since the corresponding input operator $\Phi_{\tau}:L^{\infty}([0,\tau],U)\to X_{-1}$ maps into $X$, for each $u\in L^{\infty}([0,\tau],U)$ taking values in $B_{U}$, we get the inequalities
$\int_{0}^{\tau}T_{-1}(\tau-s)b_{1}\;\mathrm{d}s\leq\Phi_{\tau}u\leq\int_{0}^{\tau}T_{-1}(\tau-s)b_{2}\;\mathrm{d}s$
in $X$ and hence the norm estimate
$\left\lVert\Phi_{\tau}u\right\rVert_{X}\leq c\max_{i=1,2}\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)b_{i}\;\mathrm{d}s\right\rVert_{X},$
where $c>0$ is such that (4.3) holds. Without loss of generality, assume that $\omega_{0}(A)<0$ and that the norm on $X_{-1}$ is given by $\left\lVert x\right\rVert_{-1}=\left\lVert(A_{-1})^{-1}x\right\rVert_{X}$ for all $x\in X_{-1}$. It follows that one can estimate $\left\lVert\Phi_{\tau}u\right\rVert_{X}$ from above by
$\left\lVert\Phi_{\tau}u\right\rVert_{X}\leq c\max_{i=1,2}\left\lVert A_{-1}\int_{0}^{\tau}T_{-1}(\tau-s)b_{i}\;\mathrm{d}s\right\rVert_{-1}=c\max_{i=1,2}\left\lVert T_{-1}(\tau)b_{i}-b_{i}\right\rVert_{-1}\to 0$
as $\tau\to 0$. In other words, $B$ is zero-class $L^{\infty}$-admissible.
(b) This now follows from the discussion at the beginning of the subsection.
(c) The arguments in the proof of (a) can be repeated mutatis mutandis. ∎
@@
Suppose that Assumption 4.1 holds and that $B$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$. Then $B$ is zero-class $C$-admissible.
Proof.
Let $\tau>0$. For each step function $u\in\mathrm{T}([0,\tau],U)$, we know that
@@
from [13, Lemma 4.3(ii)]; note that the proof of [13, Lemma 4.3(ii)] does not use any positivity properties. By assumption, there exist $b_{1},b_{2}\in X_{-1}$ such that $B(B_{U})\subseteq[b_{1},b_{2}]$. Choose $c>0$ such that (4.3) holds. If $u\in\mathrm{T}([0,\tau],U)$ with $\left\lVert u\right\rVert_{\infty}\leq 1$, then due to the positivity of the semigroup,
$\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s\right\rVert\leq c\max_{i=1,2}\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)b_{i}\;\mathrm{d}s\right\rVert=:K_{\tau}$
@@
for all step functions $u\in\mathrm{T}([0,\tau],U)$. Hence, $B$ is $\mathrm{C}$-admissible (see, [62, Remark 4.1.7] and Proposition 4.2). The zero-class $\mathrm{C}$-admissibility now follows from Theorem 4.5(c). ∎
In particular, the following generalisation of [62, Corollary 4.2.10] follows by combining Theorem 4.6 and Proposition A.2(a).
@@
Suppose that Assumption 4.1 holds and $U$ is an ordered Banach space with a unit. If $B\in\mathcal{L}(U,X_{-1})$ is positive, then it is zero-class $\mathrm{C}$-admissible.
Proposition A.2(b) gives another situation where $B$ satisfies the assumptions of Theorem 4.6. However, such a condition even yields $L^{\infty}$-admissibility (Corollary 4.11).
@@
The assumption that $U$ has a unit cannot be dropped in Corollary 4.7. Indeed, on the Banach lattice $c_{0}$, there is an example [37, Example 2.3], going back to Kato, of a positive semigroup whose generator is positive, $\mathrm{C}$-admissible, yet it is not zero-class $\mathrm{C}$-admissible.
Recall from Table 1 that an AM-space is a Banach lattice whose open unit ball is upwards directed. In particular, $c_{0}$ is an AM-space. It was shown in [62, Theorem 4.1.19] that if $U$ is an AM-space, then every positive $B\in\mathcal{L}(U,X_{-1})$ is at least $\mathrm{C}$-admissible. We generalise this in the following theorem.